OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x2 + ___x + 49

OpenStudy (cwrw238):

the two square roots involved are 2 and 7 so arithmetic on these will give you the missing number

OpenStudy (amistre64):

well, our options for this would suggest that: (2x)^2 + nx + (7)^2 comes form either: (2x+7)^2 or (2x-7)^2 ; the + on the middle term would indicate the first

OpenStudy (savvy):

since 49 is apart of the square a term must be its square root i.e. 7 and sincce the other square term is 4x^2...then the term before squaring must be its square root i.e. 2x... so (2x + 7)^2 = 4x^2 + 28x + 49

OpenStudy (anonymous):

The answer would be 4x^2+ 14x+49 You take the square root of 49 and then take that answer (which is 7) and multiply it by 2. Hope this helps!

OpenStudy (amistre64):

Zhang, that only works if the first term is an x^2; we got a 4 in the way this time to do that we would have to divide off the 4 and work it out

OpenStudy (amistre64):

but, i spose the wrong answer is the best answer :) lol

OpenStudy (amistre64):

yes, it is

OpenStudy (amistre64):

(2x+7)^2 = 4x^2 +28x + 49

OpenStudy (anonymous):

@amistre64 is my answer is wrong??

OpenStudy (amistre64):

you should be able to dbl chk it to determine that, but as is, your answer is wrong. You did not account for the 4 on the first term

OpenStudy (anonymous):

oh right lol sorry it was 28

OpenStudy (amistre64):

(ax+b)^2 = (ax)^2 + 2ab x + b^2 sqrt(b^2) * 2 = 2b but that only works if a=1, otherwise we have to adjust it for a